By volume, one alloy is 70% copper, 20% zinc, and 10% nickel.
A second alloy is 60% copper, and 40% nickel. A third alloy is
30% copper, 30% nickel, and 40% zinc. How much of each alloy
must be mixed in order to produce 1000mm3 of a final alloy that is
50% copper, 18% zinc, and 32% nickel?

Answer by stanbon(75887)

Set up three equations.
Let the three alloys by a, b, c
1st equation: Copper---: 0.7a+.6b+.3c = 0.50(1000)=500 mg
2nd equation: Zinc-----: 0.2a+ 0b+.4c = 0.18(1000)=180 mg
3rd equation: Nickel---:..............= 0.32(1000)=320 mg
---------------
This is three equations in three unknows.
I used the Matrix method on a TI calculator and found
a = 140 mg of the 1st alloy
b = 480 mg of the 2nd alloy
c = 380 mg of the 3rd alloy
==============
cheers,
Stan H.